# DataSculpt: Crafting Data Landscapes for LLM Post-Training through Multi-objective Partitioning

### v1

@@ 3.4. Phase 2: Greedy Policy
As delineated in Section 3.1, our objective encompasses augmenting the relevance (G1), homogeneity (G2), and integrity (G3) of the textual content encapsulated within each context window. This enhancement is realized through a strategic reduction in the quantity of documents concatenated into a single context sequence, amplifying the semantic congruence among these documents, and eliminating unnecessary truncations.
@@
Initially, documents within each previously formed cluster are organized in descending order by their token count. Following this organization, for each document $d_{i}$ pending allocation, the process seeks to integrate it into an existing sequence $s_{a}$ with content length that does not exceed the maximum context window capacity $L$. The sequence $s_{a}$ is selected based on achieving the highest composite score $F(d_{i},s_{a})$, which reflects a weighted consideration of semantic similarity (G1), available residual capacity (G2), and the minimization of truncation penalties (G3). Through this approach, we aim to ensure that the allocation of documents to sequences is not only computing efficient but also optimally maintains semantic consistency, maximizes capacity utilization, and preserves the integrity of individual documents within each training sequence. The strategy is pivotal in enhancing the overall effectiveness of context sequences for LLMs, by fostering a training environment that is both informationally rich and contextually coherent.
4. Experiments and Results
In this section, we describe details of our analytical study of training corpuses (Section 4.1), experimental setup (Section 4.2), the baseline methods we use for comparison (Section 4.3), and experimental results (Section 4.4).
4.1. Analytical Study on Data Corpus
The data employed for training originates from a variety of domains, including the web, books, academic papers, code repositories, etc.. The composition of these sources and their respective weight proportions are illustrated in Figure 3.
Figure 3. Domain proportion of the training corpus. We have maintained the data distribution ratios across different sources from the previous pretraining stage.
To provide further insight of our intuition, we undertook a comprehensive analysis, examining both the token quantities within individual documents among diverse domains and assessing the effectiveness of semantic clustering across different methodologies.
4.1.1. Length Analysis
We conducted a statistical analysis of the token number characteristics across various domains, as illustrated in Table 1. We also categorize the tokenized documents into different length intervals based on token quantities across each domain. This process enables us to derive the distribution of token lengths for individual documents situated within specific domains. The proportional distributions across domains are presented as a normalized stacked horizontal bar chart in Figure 1, which provides a hierarchical view of the domains according to their prevalence of longer texts. It reveals that data originating from the book and academic paper domains predominantly consist of long documents, with more than half of the data exceeding sequence lengths of 64K tokens. In contrast, academic papers tend to feature medium-length documents (ranging from 4K to 32K tokens), while web-sourced data are characterized by a preponderance of short documents, almost exclusively comprising texts shorter than 4K tokens. Consequently, it is evident that web data does not inherently lend itself to extending model training windows due to its shorter length, necessitating the concatenation of multiple individual documents to form longer sequences suitable for training.
Min Len 667 322 127 173 696 289 132 23
Avg Len 176,156 216,969 20,555 2,578 6,216 886 358 224
Table 1. Token statistics among domains. We conducted an analysis of the maximum, minimum, and average token counts from different domains following tokenization. Data sourced from the web tends to have a shorter average length, necessitating the concatenation of multiple documents to achieve extended sequences suitable for long-context training.
The attribute that individual documents originating from books and academic papers tend to have a longer average length significantly increases their propensity to occupy the entirety of a context sequence when training models under expanded contexts, such as 32K or 64K tokens. Conversely, web-derived data predominantly exhibits a shorter average document length compared to other domains, necessitating the concatenation of multiple documents to form sequences of adequate length during the training process of extended sequence lengths.
This observation underscores the intrinsic variability in document length across different domains. However, a balanced domain mixture is important for extend context training, where experiments conducted by Fu et al. (Fu et al., 2024) demonstrated the importance of upsampling longer sequences while preserving the domain mixture within the pretraining corpora for effective context scaling. Therefore, it is necessary to adopt flexible data organization strategies tailored to accommodate the varying length distributions across domains. Specifically, for domains characterized by longer documents, such as books and academic papers, leveraging the natural length of these documents can optimize the utilization of context windows during training. On the other hand, for web-sourced data, where a majority of the individual documents are insufficient to span the entire context sequence, a strategy involving the concatenation of multiple documents to construct longer sequences becomes imperative. This insight holds substantial significance for our work. The approach ensures that the training process effectively captures the broader context as well as takes into account the characteristics of different data sources, thereby enhancing the model’s overall ability to understand and generate coherent text across varying lengths and domains.
Considering the distribution of document lengths and proportion of each domain, we have strategically organized the web-produced English data within our training corpus, which constitutes 30.5% of the total training data.
4.1.2. Semantic Clustering Analysis
@@
In evaluating the effectiveness of semantic clustering, our analysis encompasses two key dimensions: the statistical features related to the document quantities within each cluster, and the application of t-SNE (Van der Maaten and Hinton, 2008) for dimensionality reduction to project and analyze clusters. These clusters were derived from two distinct approaches: the greedy traversal of incomplete graphs as described in ICLM (Shi et al., 2023), for which we refer to the official implementation  [^1] https://github.com/swj0419/in-context-pretraining, and the semantic clustering achieved through an ISODATA variant (Algorithm 1) implemented by our DataSculpt.
Numerical Narratives.  Initially, we performed a quantitative analysis of the number of clusters generated by two distinct methodologies, as detailed in Table 2. It revealed that the clusters produced through greedy graph traversal in ICLM totaled 323,428,858, markedly surpassing the 611,981 clusters formed via semantic clustering with Algorithm 1, which suggests a more fragmented clustering approach by the former method. Moreover, clusters initiated at the commencement of the traversal process encompassed a substantially larger number of documents compared to those established later, leading to a pronounced imbalance in document distribution. Specifically, a significant skew was observed in the distribution of clusters from ICLM, with 53.56% comprising isolated single-document clusters. The dominance of such minimal-scale clusters hinders the ability to adequately fill the context window during the concatenation process for long sequences, thereby inadequately mitigating cross attention in training despite the potential for high semantic similarity among documents. Conversely, DataSculpt exhibited a more equitable distribution of cluster sizes throughout the semantic clustering process, enabling the majority of clusters containing semantically akin documents to extend to the length required for a long context sequence. This strategy efficiently capitalizes on the advantages of clustering similar documents, which significantly enhances the model’s training efficiency by promoting semantic coherence within clusters.
T-SNE Visualization Journeys.  Data visualization is crucial for helping decision-making during massive data management (Qin et al., 2019). Here we also employed the dimensionality reduction technique t-SNE to visualize both the semantic clusters generated via Algorithm 1 and those identified through greedy graph traversal mechanism on an incomplete graph as described in ICLM. For the latter approach, each document was connected to its 20 nearest neighbors, facilitating the construction of the incomplete graph.
@@ 4.2. Experimental Setup
We continue pretraining a Transformer decoder-only 7B model (Dong et al., 2024) similar to LLaMA framework (Touvron et al., 2023) by different sources with the same distribution ratios to the previous pretraining stage, where we structured the data derived from the web (English) domain in accordance with a previously defined multi-objective partition allocation strategy (Algorithm 2). The details of training datasets can be found in Section 4.1.
Data Processing Details.  During the data preprocessing phase, we tokenize raw documents from the training corpus (Baichuan, 2023), which collectively encompass 855 billion tokens, and then truncated documents at varying sequence lengths of 16K, 32K, and 64K. In conducting vector embedding across our training datasets, we employed the BGE-m3 model (Chen et al., 2024b) as our embedding model, encoding each document into a $1024$-dimensional vector. The encoding process necessitated approximately 9 days using a 1,044,090 GPU cluster. Following this, we constructed an HNSW index based on FAISS index (Johnson et al., 2019) to facilitate efficient similarity search for semantic clustering (Section 3.3). The greedy semantic-driven partition algorithm (Section 3.4) within each cluster was executed on a 4114 CPU Ray cluster.
Model Details.  The model comprises a Transformer decoder with 32 attention heads across 32 layers. It features a hidden layer size of 4096, coupled with a dimensionality of 11008 for the forward propagation layers, and the base of Rotary Positional Encoding (RoPE) (Su et al., 2024) is modified to accommodate longer context. We train with 16K, 32K, 64K sequence lengths using 15B tokens for the 7B base model (recall that base model has the 16K context length), applied on 112 H800 GPUs across 14 nodes. We employ global batch sizes of 224, 112, and 56 sequences per step for context lengths of 16K, 32K, and 64K, which maintains a uniform number of tokens for model parameter updates. The learning rate is set to 3e-4 with linear warm-up over the first 1,000 optimization steps and cosine decay schedule.
@@ 4.4.1. General Understanding
General understanding tasks require the model to comprehend the text holistically, extracting key elements and answering queries, which focuses on broad understanding over specific task execution. We evaluate the 7B natural language models with zero-shot on MMLU (Hendrycks et al., 2021b, a), TyDiQA (Clark et al., 2020), BBH (Srivastava et al., 2022) and few-shot in-context learning on TREC (Li and Roth, 2002; Hovy et al., 2001). MMLU encompasses 57 distinct tasks featuring multiple-choice questions from diverse knowledge domains such as humanities, social sciences and hard sciences. TyDiQA is a multilingual question answer dataset containing 204K pairs across 11 topologically diverse languages, which captures language phenomena absent in English-only corpora. BBH focuses on a suite of 23 challenging tasks from BIG-Bench that require multi-step reasoning, where previous language model evaluations fell short of average human-rater performance. TREC is a text classification challenge aimed at categorizing queries into predefined classes, thereby assessing a model’s ability to accurately classify questions, facilitating the determination of suitable responses or actions. Table 3 contains more details about these tasks. Each dataset includes multiple subtasks, and each subtask comes with few-shot examples.
Dataset # Tasks (Labels) # Instances Answer Type
MMLU 57 14,079 Letter options
TyDiQA 9 1,713 Span
BBH 23 920 COT and answer
TREC 6(1st)-47(2nd) 500 Letter options
Table 3. Statistics of evaluation datasets for general understanding. The evaluation tasks cover different kinds of types, including question answering and text classification.
In Table 4, we compare our models trained via DataSculpt with other baseline models in general understanding across various context window size. DataSculpt not only aligns but also slightly outperforms the baselines, particularly in areas requiring multi-hop logical reasoning tasks (+4.88% compared to random sampling, +4.38% compared to ICLM).
@@
4.4.2. Retrieval Augmentation
Retrieval augmentation involves LLMs to access external knowledge sources, such as databases or text corpora to retrieve pertinent information as required, effectively extending their memory and knowledge capabilities beyond what is available in their pretrained state. In assessing the performance of models on retrieval augmentation tasks, we focus on the models’ synthetic key-value retrieval performance (Liu et al., 2024) under settings of 140 and 300 key-value pairs and conduct a zero-shot evaluation using the TriviaQA dataset (Joshi et al., 2017). Further experiments are also undertaken on the “needle in a haystack” task (Kamradt, 2023; Fu et al., 2024) with a maximum of 64K tokens to gain a comprehensive understanding of the effectiveness of these models in identifying specific information within extensive contexts. TriviaQA comprises over 650K question-answer-evidence triples and 95K question-answer pairs labeled with evidence passages curated by trivia enthusiasts. The key-value retrieval task (Liu et al., 2024) has evolved into a standard diagnostic tool for exploring contextual capabilities, where models are presented with a set of key-value pairs and tasked to retrieve a value corresponding to a specific key from a set of key-value pairs. The “needle in a haystack” task is designed to evaluate the models’ capacity for in-context retrieval within extended contexts, further testing their ability to navigate and extract pertinent information from dense information sets.
@@ 4.4.3. Summarization
Summarization tasks involve generation of concise and coherent abstracts from a given source text, which evaluate a model’s capacity to distill key information, maintain the original text’s intent, and present it in a brief and comprehensible manner. Compared to traditional QA and text classification tasks, which frequently rely on local context for solutions, summarization requires a broader understanding of the entire document (Bai et al., 2023). This capability is more closely associated with LLMs’ long-context capabilities. We assess the post-trained models in a zero-shot setting on GovReport (Huang et al., 2021), QMSum (Zhong et al., 2021), and employ few-shot in-context learning on SAMSum (Gliwa et al., 2019). GovReport is a large-scale dataset for long document summarization, featuring approximately 19.5K U.S. government reports paired with expert-crafted abstractive summaries. QMSum is a benchmark for query-based multi-domain meeting summarization, encompassing annotated query-summary pairs spanning 232 meetings in diverse domains including product development, academia, and committee discussions. SAMSum is specifically curated for dialogue summarization, comprising messenger-style conversations accompanied by human-annotated summaries. Additional details about these tasks can be found in Table 7. Here we utilize the ROUGE-L (Lin, 2004) metric, an N-gram based measure extensively adopted in QA and summarization tasks. Improvements in summarization tasks achieved through DataSculpt are documented in Table 8, showcasing enhancements of 21.21% over random sampling and 18.18% over ICLM.
Dataset Source Avg Len Query Metric
GovReport Government report 8,734 $\times$ Rouge-L
QMSum Meeting 10,614
SAMSum Dialogue 6,258 $\times$
Table 7. Statistics of evaluation datasets for summarization. “Source” refers to the provenance of the dataset, and “Avg Len” is calculated based on word counts. “Query” signifies whether the dataset conforms to the query-based format.
@@ 4.4.4. Reading Comprehension
Dataset Avg Len Classification Metric
NarrativeQA 18,409 Single-Doc F1
Qasper 3,619
HotPotQA 9,151 Multi-Doc
2WikiMultihopQA 4,887
Table 9. Statistics of evaluation datasets for reading comprehension. “Avg Len” is calculated based on word counts.
@@
Reading comprehension assesses the capacity of models to generate responses to queries, grounded in the context provided within a given textual passage. We evaluate the models’ ability in reading comprehension with zero-shot on NarrativeQA (Kočiskỳ et al., 2018), Qasper (Dasigi et al., 2021), HotPotQA (Yang et al., 2018) and 2WikiMultihopQA (Ho et al., 2020). NarrativeQA consists of long stories in the form of books and movie scripts, accompanied by human-crafted questions designed to assess reading comprehension. Qasper is a dataset comprising 5,049 questions derived from 1,585 NLP papers. The questions are formulated by NLP experts based solely on paper titles and abstracts, while answers are addressed by a different group of NLP practitioners. HotpotQA involves 113,000 two-hop question-answer pairs with four key features, authored directly by native speakers given two related paragraphs. 2WikiMultihopQA is a multi-hop dataset that has comprehensive explanations from question to answer derived from Wikipedia and Wikidata, synthesized with manually designed templates to prevent shortcut-based solutions. NarrativeQA and Qasper fall into the category of single-document question-answering (QA) datasets, whereas HotPotQA and 2WikiMultihopQA are recognized as Wikipedia-based, multi-hop QA datasets. Evaluation across these datasets is conducted using the F1 metric, as outlined in Table 9. Enhancements attributed to DataSculpt in reading comprehension tasks are detailed in Table 10 and Table 11.
@@ 4.4.5. Code Completion
Code completion serves as a critical and challenging task for evaluating a model’s long context modeling ability, which facilitates user coding by predicting and completing code segments based on previous inputs and contextual understanding. For code completion tasks, we assess the models using zero-shot learning on the HumanEval dataset (Chen et al., 2021), LCC (Guo et al., 2023), and RepoBench-P (Liu et al., 2023), where the latter two datasets are specifically designed for long code modeling evaluation. HumanEval contains 164 handwritten programming problems, each accompanied by a function signature, a documentation string, the main body of the code, and an average of 7.7 unit tests per challenge to assess correctness. LCC is developed by selecting single-file code snippets from GitHub based on length criteria. RepoBench-P is curated from GitHub repositories, integrating retrieval and code completion tasks into a cohesive pipeline, where the goal is to first identify and retrieve the most pertinent code snippets based on the context within a file, and then predict the most appropriate subsequent line of code based on the retrieved snippets as well as the in-file context. Further information regarding these tasks is provided in Table 12, where we applied execution-based accuracy (Pass@1) for HumanEval evaluation and Edit Sim (Levenshtein distance) for the remaining two code generation tasks.
Dataset Language Level Metric
HumanEval Python function Pass@1
LCC Python/Java/C# file Edit Sim
RepoBench-P Python/Java repository
Table 12. Statistics of datasets for code completion. “Language” refers to the programming languages included within the dataset, and “Level” denotes the task’s focus granularity.

### v2

@@ 15      $r(s_{a^{*}})\leftarrow r(s_{a^{*}})-l(d_{i})$;
As delineated in Section 3.3, our objective encompasses augmenting the relevance (G1) and homogeneity (G2), while maintaining integrity (G3) of the textual content encapsulated within each context window. This enhancement is realized through a strategic reduction in the quantity of documents concatenated into a single context sequence, amplifying the semantic congruence among these documents, and eliminating unnecessary truncations.
@@
Initially, documents within each previously formed cluster are organized in descending order by their token count (Line 2). Following this organization, for each document $d_{i}$ pending allocation, the process seeks to integrate it into an existing sequence $s_{a}$ with content length that does not exceed the maximum context window capacity $L$ (Line 3-14). The sequence $s_{a}$ is selected based on achieving the highest composite score $F(d_{i},s_{a})$, which reflects a weighted consideration of semantic similarity (G1, Line 8), available residual capacity (G2, Line 9), and the minimization of truncation penalties (G3, Line 10). Through this approach, we aim to ensure that the allocation of documents to sequences is not only computing efficient but also optimally maintains semantic consistency, maximizes capacity utilization, and preserves the integrity of individual documents within each training sequence (Line 11-12). The strategy is pivotal in enhancing the overall effectiveness of context sequences for LLMs, by fostering a training environment that is both informationally rich and contextually coherent.
4.3. Semantic Clustering Analysis
@@
To demonstrate the robustness of our approach, we compared the effectiveness of semantic clustering across different methodologies. Our analysis encompasses two key dimensions: the statistical features related to the document quantities within each cluster, and the application of t-SNE (Van der Maaten and Hinton, 2008) for dimensionality reduction to project and analyze clusters. These clusters were derived from two distinct approaches: the greedy traversal of incomplete graphs as described in ICLM (Shi et al., 2023), for which we refer to the official implementation  [^1] https://github.com/swj0419/in-context-pretraining, and the semantic clustering achieved through an ISODATA variant (Algorithm 1) implemented by our DataSculpt.
4.3.1. Numerical Narratives
Initially, we performed a quantitative analysis of the number of clusters generated by two distinct methodologies, as detailed in Table 2. It revealed that the clusters produced through greedy graph traversal in ICLM totaled 323,428,858, markedly surpassing the 611,981 clusters formed via semantic clustering with Algorithm 1, which suggests a more fragmented clustering approach by the former method. Moreover, clusters initiated at the commencement of the traversal process encompassed a substantially larger number of documents compared to those established later, leading to a pronounced imbalance in document distribution. Specifically, a significant skew was observed in the distribution of clusters from ICLM, with 53.56% comprising isolated single-document clusters. The dominance of such minimal-scale clusters hinders the ability to adequately fill the context window during the concatenation process for long sequences, thereby inadequately mitigating cross attention in training despite the potential for high semantic similarity among documents. Conversely, DataSculpt exhibited a more equitable distribution of cluster sizes throughout the semantic clustering process, enabling the majority of clusters containing semantically akin documents to extend to the length required for a long context sequence. This strategy efficiently capitalizes on the advantages of clustering similar documents, which significantly enhances the model’s training efficiency by promoting semantic coherence within clusters.
4.3.2. T-SNE Visualization Journeys
Data visualization is crucial for helping decision-making during massive data management (Qin et al., 2019). Here we employed the dimensionality reduction technique t-SNE to visualize the semantic clusters generated via Algorithm 1 and those identified through greedy graph traversal on an incomplete graph as described in ICLM. For the latter approach, each document was connected to its 20 nearest neighbors, facilitating the construction of the incomplete graph.
@@ 4.3.2. T-SNE Visualization Journeys
5. Experiments and Results
In this section, we describe details of our experimental setup (Section 5.1), the baseline methods we use for comparison (Section 5.2), and experimental results (Section 5.3).
@@ 5.1. Experimental Setup
We continue pretraining a Transformer decoder-only 7B model (Dong et al., 2024) similar to LLaMA framework (Touvron et al., 2023) by different sources with the same distribution ratios to the previous pretraining stage, where we structured the data derived from the web (English) domain in accordance with a previously defined multi-objective partition allocation strategy (Section 4). The details of training datasets can be found in Section 4.3.
Data Processing Details.  During the data preprocessing phase, we tokenize raw documents from the training corpus (Baichuan, 2023), which collectively encompass 855 billion tokens, and then truncated the documents at varying sequence lengths of 16K, 32K, and 64K. In conducting vector embedding across our training datasets, we employed the BGE-m3 model (Chen et al., 2024b) as our embedding model, encoding each document into a $1024$-dimensional vector. The encoding process necessitated approximately 9 days using a 1,044,090 GPU cluster. Following this, we constructed an HNSW index based on FAISS index (Johnson et al., 2019) to facilitate efficient similarity search for semantic clustering (Section 4.1). The greedy semantic-driven largest-fit partition algorithm (Section 4.2) within each cluster was executed on a 4114 CPU Ray cluster.
Model Details.  The model comprises a Transformer decoder with 32 attention heads across 32 layers. It features a hidden layer size of 4096, coupled with a dimensionality of 11008 for the forward propagation layers, and the base of Rotary Positional Encoding (RoPE) (Su et al., 2024) is modified to accommodate longer context. We train with 16K, 32K, 64K sequence lengths using 15B tokens for the 7B base model (recall that base model has the 16K context length), applied on 112 H800 GPUs across 14 nodes. We employ global batch sizes of 224, 112, and 56 sequences per step for context lengths of 16K, 32K, and 64K, which aims at maintaining a uniform number of tokens for model parameter updates. The learning rate is set to 3e-4 with linear warm-up over the first 1,000 optimization steps and cosine decay schedule.
@@ 5.3.1. General Understanding
General understanding tasks require the model to comprehend the text holistically, extracting key elements and answering queries, which focuses on broad understanding over specific task execution. We evaluate the 7B continual pretrained natural language models with zero-shot on MMLU (Hendrycks et al., 2021), TyDiQA (Clark et al., 2020), BBH (Srivastava et al., 2022) and few-shot in-context learning on TREC (Li and Roth, 2002). The evaluation tasks cover different kinds of types, including question answering and text classification. Each dataset includes multiple subtasks, and each subtask comes with few-shot examples. In Table 3, we compare our models trained via DataSculpt with other baseline models in general understanding across various context window size. DataSculpt not only aligns but also slightly outperforms the baselines, particularly in areas requiring multi-hop logical reasoning tasks (+4.88% compared to random sampling, +4.38% compared to ICLM).
@@
5.3.2. Retrieval Augmentation
Retrieval augmentation involves LLMs to access external knowledge sources, such as databases or text corpora to retrieve pertinent information as required, effectively extending their memory and knowledge capabilities beyond what is available in their pretrained state. In assessing the performance of models on retrieval augmentation tasks, we focus on the models’ synthetic key-value retrieval performance (Liu et al., 2024) under settings of 140 and 300 key-value pairs and conduct a zero-shot evaluation using the TriviaQA dataset (Joshi et al., 2017). Further experiments are also undertaken on the “needle in a haystack” task (Kamradt, 2023; Fu et al., 2024) with a maximum of 64K tokens to gain a comprehensive understanding of the effectiveness of these models in identifying specific information within extensive contexts.
@@ 5.3.3. Summarization
Summarization tasks involve generation of concise and coherent abstracts from a given source text, which serves as a measure of a model’s ability to distill key information, maintain the original text’s intent, and present it in a brief and comprehensible manner. Compared to traditional QA and text classification tasks, which predominantly depend on local context for deriving solutions, summarization requires a broader understanding of the entire document (Bai et al., 2023). This capability is more closely associated with the long-context processing capabilities of LLMs.
We assess the continued pretrained models in a zero-shot setting on real-world benchmarks including GovReport (Huang et al., 2021) as well as QMSum (Zhong et al., 2021), and employ few-shot in-context learning on SAMSum (Gliwa et al., 2019). Here we utilize the ROUGE-L (Lin, 2004) metric, an N-gram based measure extensively adopted in QA and summarization tasks. Improvements in summarization tasks achieved through DataSculpt are documented in Table 6, showcasing enhancements of 21.21% over random sampling and 18.18% over ICLM.
@@ 5.3.4. Reading Comprehension
Reading comprehension assesses the capacity of models to generate responses to queries, grounded in the context provided within a given textual passage. We evaluate the models’ ability in reading comprehension with zero-shot on NarrativeQA (Kočiskỳ et al., 2018), Qasper (Dasigi et al., 2021), HotPotQA (Yang et al., 2018) and 2WikiMultihopQA (Ho et al., 2020). NarrativeQA and Qasper fall into the category of single-document question-answering (QA) datasets, whereas HotPotQA and 2WikiMultihopQA are recognized as Wikipedia-based, multi-hop QA datasets. Evaluation across these datasets is conducted using the F1 metric, and enhancements attributed to our DataSculpt in reading comprehension tasks are further detailed in Table 7 and Table 8.
@@ 5.3.5. Code Completion
Code completion serves as a critical and challenging task for evaluating a model’s long context modeling ability, which facilitates user coding by predicting and completing code segments based on previous inputs and contextual understanding. For code completion tasks, we assess the models using zero-shot learning on the HumanEval dataset (Chen et al., 2021), LCC (Guo et al., 2023), and RepoBench-P (Liu et al., 2023), where the latter two datasets are specifically designed for long code modeling evaluation. Here we applied execution-based accuracy (Pass@1) for HumanEval evaluation and Edit Sim (Levenshtein distance) for the remaining two code generation tasks.
@@
6. Related Work
6.1. Data-Centric Long-Context Training
Training Large Language Models (LLMs) with expanded context windows facilitates the execution of tasks that surpass the capabilities defined by the existing paradigm, such as multi-document question answering (Caciularu et al., 2023), long document summarization (Bai et al., 2023), and enhanced retrieval functionalities (Liu et al., 2024). However, LLMs’ efficacy in handling long context does not consistently meet the expected standards. Data-focused approaches hold substantial importance (Wang et al., 2023) for enhancing LLMs’ long-context capabilities.
The construction of long-context data involves more than a mere concatenation of short documents, which can lead to cross-document attention within a sequence. In these circumstances, formulating high-quality or domain-specific long sequences holds significant research value (de Vries, 2023). There have been data-centric efforts made in improving LLMs’ capabilities for long context processing. ICLM (Shi et al., 2023) and SPLICE (Staniszewski et al., 2023) combine similar training instances through a sophisticated retrieval process, which emphasizes the critical role of high-quality long-context training data in enhancing the context expansion capabilities of LLMs without necessitating extensive data resources. Additionally, Krell et al. (Krell et al., 2021) deliberate on the challenges associated with the conventional concat-and-chunk approach, proposing an approximate bin-packing algorithm as a potential solution. As an improvement, dataset decomposition (Pouransari et al., 2024) and Best-fit Packing (Ding et al., 2024) both propose the sequence length aware optimization technique to minimize truncations, thereby facilitating more efficient pretraining on long sequences and achieving scalability relative to the size of dataset. Moreover, ProLong (Chen et al., 2024a) introduces an evaluation framework for assessing the quality of long-context data, further enriching the field’s methodological toolkit. Nonetheless, these methods have not performed a systematic framework tailored for data construction in long context training. Our work conceptualizes data management for extended-context training as a multi-objective combinatorial optimization problem, and provide a comprehensive framework for long sequence data organization.
6.2. MOCO-Augmented LLMs
Existing works have integrated Multi-objective combinatorial optimization (MOCO) problems with relevant concerns in the LLMs field. Wang et al. (Wang et al., 2024) employed the MOCO approach within the Reinforcement Learning from Human Feedback (RLHF) framework for LLMs, where the reward models aggregate multiple reward objectives with coefficients into a scalar score. Akiba et al. (Akiba et al., 2024) viewed model merging as a MOCO challenge, and introduced an innovative approach that leverages evolutionary algorithms to facilitate the automated development of robust foundation models. Furthermore, integration of MOCO with data-centric LLMs have been explored for training data selection (Du et al., 2023; Xu et al., 2023; Chen et al., 2024a), which applies heuristic techniques to assign weights to different data evaluation metrics and facilitates a phased resolution approach. Considering the extensive scale of training dataset, data construction for long context is treated as a MOCO challenge, for which the coarse-to-fine algorithm is identified as a suitable resolution strategy.